A tennis ball is served and travels the length of court L in time T. a) Write an equation for the balls average horizontal speed. b) Calculate the average speed of a ball travelling 24.0m across the court in 0.60s
Answer #1

a)
Average speed = distance travelled/time taken

= L/T

b) V_avg = L/T

= 24/0.6

= 40 m/s